OpenStudy (anonymous):

Suppose f(x)dx = 8, (f(x)dx)^2=12, g(x)dx=2, (g(x)dx)^2=3. Evaluate cf(z)dz. All of these have integral of a to b.

OpenStudy (aum):

\[ \int_a^bf(x)dx = 8 \\ \int_a^bcf(z)dz =c\int_a^bf(z)dz = c * 8 = 8c \]